OpenStudy (anonymous):

prove : 1+tan^2 A / 1-tan^2 A = sec^2 A

OpenStudy (lalaly):

are u sure its right?

OpenStudy (anonymous):

yes its written like that.

OpenStudy (anonymous):

\[(1+\tan^2 A)/(1-\tan^2 A)\]

OpenStudy (lalaly):

oh lols ok wait

OpenStudy (anonymous):

i really appreciate your help lalaly.thank you so much.take your time.

OpenStudy (lalaly):

\[\sec^2A= 1+\tan^2A\] \[\tan^2A = \sec^2A-1\] \[\frac{1+\tan^2A}{1-\tan^2A} = \frac{\sec^2A}{1-(\sec^2A-1)} = \frac{\sec^2A}{\sec^2A} = 1\] i am sure there is sth wrong here i checked wolfram too http://www.wolframalpha.com/input/?i=%281%2Btan^2+A%29+%2F+%281-tan^2+A%29 check ur queation again

OpenStudy (lalaly):

you hav to prove it is equal to sec(2A) not sec^2A

OpenStudy (lalaly):

1 +tan ² x = ----------------- 1 - tan ² x 1 + sin ² x / cos ² x = -------------------------- 1 -sin ² x / cos ² x cos ² x +sin ² x = ------------------- cos ² x - sin ² x 1 = ------------- cos ² x sin ² x 1 = ---------- cos(2x) = sec(2x)

OpenStudy (anonymous):

oh got it man.thanks :) sorry for the late response btw . i just took a shower.

OpenStudy (anonymous):

thanks again dude :)